Dolores used to work as a high school teacher for $55,000 per year but quit in order to start her own catering business. To buy the necessary equipment, she withdrew $30,000 from her savings (which paid 2 percent interest per year) and borrowed $30,000 from her uncle, to whom she pays 3 percent interest per year. Last year she paid $25,000 for ingredients and had revenue of $75,000. She asked Louis, an accountant, and Greg, an economist, to calculate her profit for her. What did they say?

Question 1 options:

Louis said her profit was $49,100, and Greg said she lost $6500.

Louis said her profit was $49,100, and Greg said her profit was $6500.

Louis said her profit was $50,000, and Greg said she lost $5000.

Louis said her profit was $4500, and Greg said her profit was $33,500.

The Wheeler Wheat Farm sells wheat to a grain broker in Regina, Saskatchewan. Since the market for wheat is generally considered to be competitive, what would the Wheeler Wheat Farm choose to do to maximize its profit?

Question 2 options:

It would produce the quantity at which average total cost is minimized.

It would produce the quantity at which the average fixed cost is minimized.

It would sell its wheat at a price where marginal cost is equal to average total cost.

It would produce the quantity at which market price is equal to the farm’s marginal cost of production.

A profit-maximizing firm in a competitive market is able to sell its product for $8. At its current level of output the firm’s average total cost is $11. Its marginal-cost curve crosses the marginal revenue curve at an output level of 10 units. At that point, what does the firm experience?

Question 3 options:

a loss of more than $30

a loss of exactly $30

a profit of exactly $30

a profit of more than $30

When the entry and exit behaviour of firms in an industry does not affect a firm’s cost structure, what is the shape of the long-run market supply curve?

Question 4 options:

It must be horizontal.

It must be upward.

It is downward sloping then upward sloping.

It is upward sloping then downward sloping.
Answer #1

Answer 1. Louis said her profit was $49,100, and Greg said she lost $6500

Reason- Loss of interest rate= 2%*30000= 600

Interest paid to uncle= 3%*30000= 900

Revenue= 75000

Cost= 25000

Profit according to Louis-

75000-25000-900= $49100

Profit According to Greg= $49100- 55000-600= -$6500

Answer 2. It would produce the quantity at which market price is equal to the farm’s marginal cost of production.

Reason- Under perfect comeption firm produces where P=MC. P=MR as firms are price taker and cannot control the prices.

Answer 3. A loss of exactly $30

Reason- Profit= (P-ATC)*Q= (8-11)*10= -$30

Loss=$30

Answer 4. It must be horizontal

Reason- In the long run supply curve will be perfectly elastic when the entry or exit behaviour of firms in industry does not affect a firm's cost structure.

If it helps kindly upvote